From the set {45, 90, 180}, use substitution to determine which value of x makes the equation true.
1,260 ÷ x = 28
A.
45
B.
none of these
C.
90
D.
180

Answer by Alan3354(69443)

1,260 ÷ x = 28
--------
1260/x = 28
1260 = 28x
x = 1260/28
x = 45